OpenStudy (adobohobo):

find the standard form for the equation of the line which passes through the point (-1,-2) and is parallel to the line that has an equation of 6x+2y=4

OpenStudy (hey199):

j

OpenStudy (adobohobo):

what? @Hey199

OpenStudy (adobohobo):

i tried to follow what this video did but i ended up with y=3x-4 but that isn't the correct answer

OpenStudy (hey199):

oh my bad didnt meen to say j

OpenStudy (adobohobo):

a 3x+y=5 b 3x+y=-5 c 3x+y=-1 d 3x+y=-7 these are the options

OpenStudy (perl):

6x+2y = 4 Did you solve for y first? I suggest that since we need the slope.

OpenStudy (adobohobo):

like as in y=2-3x? @perl

OpenStudy (hey199):

y = -3x +10

OpenStudy (adobohobo):

shoot what am i doing wrong

OpenStudy (hey199):

no

OpenStudy (adobohobo):

sorry for being slow rn its 247 am lol

OpenStudy (adobohobo):

where i am at

OpenStudy (perl):

Yes that's correct. y = -3x + 2 Parallel lines have the same slope so the slope of your new line will be -3. Can you finish it from here?

OpenStudy (adobohobo):

@perl would 3x+y=5 be the answer

OpenStudy (retireed):

Can you show me your work. so I can check your process?

OpenStudy (retireed):

I want to see if you worked it out or just do a guess.